OpenStudy (anonymous):

Answer Please ? Ill Medal Which algebraic expression is equivalent to the expression below? 6(4x + 3) + 5x A. 29x + 3 B. 24x + 18 C. 29x + 18 D. 9x + 3

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

B ?

OpenStudy (skullpatrol):

why?

OpenStudy (skullpatrol):

don't guess :)

OpenStudy (anonymous):

I didnt :/ ... I multiplied 6 by 4 and 6 by 3

OpenStudy (anonymous):

it would be C Because 6x*4(x)+5x=29x And 6*3=18

OpenStudy (anonymous):

Its C

OpenStudy (aaronandyson):

? 6(4x + 3) + 5x Distribute 6 to 4x ans 3 (6*4x + 3*6) + 5x (24x + 18) + 5x 24x+18+5x 24x+5x+18 = ?

OpenStudy (anonymous):

kk thanks

OpenStudy (aaronandyson):

So,what do you think the answer is ?

OpenStudy (anonymous):

she already got the answer it was C